# Hierarchical Balance Packing: Towards Efficient Supervised Fine-tuning for Long-Context LLM

### v1

@@ 4.2 Balance Packing
| HBP | 72.0 | 76.5 | 83.1 | 72.6 | 41.4 | 89.9 | 78.1 | 87.3 - | 50.3 | - | 23.8 (2.4) |
4.3 Dynamic Training Pipeline
Since HBP involves multi-level inputs, it is essential to design a dynamic training pipeline, enabling hot switching of different packing groups. It incorporates an adaptive sequential parallel, a curriculum learning strategy, and a stable loss normalizer.
Adaptive Sequential Parallel: Each packing group is equipped with an optimal training strategy . We use alternative training between packing groups with the best SP degree and GC configuration.
Curriculum Learning Strategy: Training on long-context tasks presents challenges because initiating training without instructional capabilities can result in significant fluctuations in the training loss, as illustrated in Figure 3. Thanks to our inherent hierarchical structure, it is much easier to adopt a curriculum learning strategy that starts with general short-context data during the early stages of training. As the training progresses, we shift to a hybrid approach that combines both general short- and long-context data.
Stable Loss Normalizer: The training stability introduced by data packing is an important problem as it impacts the data distribution. Previous work (Bai et al., 2024a) has analyzed loss calculation, identifying two primary loss normalizers Token Mean and Sample Mean:
where represents the batch size, and denotes the number of loss tokens in the current batch. They argued that instability lies in the inconsistent loss of normalization. (Lambert et al., 2024b) also proposes sum loss without normalization to mitigate the effect of norm discrepancies. However, the sum loss introduces a trade-off: as the sequence length increases, the gradient values escalate disproportionately (1e+5), as shown in the left part of Figure 4.
To address the above issues, we empirically find (Average Token) of the global batch size could served as a Stable Loss Normalizer:
where represents the number of loss tokens that need to be calculated for the current batch .
5 Experiments
5.1 Experimental Setup
We use large-scale datasets: Tulu3 (32K) (Lambert et al., 2024a) (general task), and LongCite (128K) (Zhang et al., 2024) (long context task). These datasets have proven effective in previous research, significantly enhancing model performance across knowledge, reasoning, mathematics, coding, and instruction-following. Moreover, they improve the model’s ability to handle varying lengths, from 0.1K to 128K tokens.
Implementation Details. We conduct experiments with the following models: LLaMA 3.1 (Dubey et al., 2024), Qwen-2.5 (Yang et al., 2024), and DeepSeek-V2 (236B). Most models are trained on 32x H100 80GB GPUs using the DeepSpeed (Rajbhandari et al., 2020), while DeepSeek-V2 (236B) is trained with the Megatron-LM (Shoeybi et al., 2019) with 256x H100 80G GPUs. We conducted ablation experiments using the LLaMA3.1-8B model. For the Longsite dataset, approximately 2k samples are uniformly sampled. The loss normalizer for baselines without special instructions is Token-Mean, while HBP uses Ave-Token. GPU days are the evaluation metric to estimate the total training time.
@@ 5.1 Experimental Setup
| Tulu3 + Longcite | 128K | 56.0 | 44.0 | 67.5 |
| Model | Hierarchical | Balance | ABR | CR | AVE | GPU Days (speed up) |
| LLaMA3.1-8B | 0.506 | 1.0 | 56.0 | 5.22 (1.0) | ||
| LLaMA3.1-8B | ✓ | 0.288 | 0.173 | 56.4 | 4.51 (1.2) | |
| LLaMA3.1-8B | ✓ | ✓ | 0.002 | 0.173 | 56.6 | 3.73 (1.40) |
| LLaMA3.1-70B | 0.506 | 1.0 | 72.2 | 44.4 (1.0) | ||
| LLaMA3.1-70B | ✓ | 0.288 | 0.173 | 72.8 | 33.3 (1.25) | |
| LLaMA3.1-70B | ✓ | ✓ | 0.002 | 0.173 | 72.2 | 31.1 (1.43) |
Evaluation. We conducted a comprehensive evaluation of the LLM’s performance using OpenCompass (Contributors, 2023a). For general tasks, several benchmark datasets were assessed, including MMLU (Hendrycks et al., 2021), MMLU PRO (Wang et al., 2024b), CMMLU (Li et al., 2023), BBH (Suzgun et al., 2022), Math (Saxton & Kohli, 2019), GPQA Diamond (Rein et al., 2024), GSM8K (Cobbe et al., 2021), HellaSwag (Zellers et al., 2019), MathBench (Liu et al., 2024), HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), IFEval (Zhou et al., 2023), and Drop (Dua et al., 2019). For long-context tasks, the evaluation including Ruler (Hsieh et al., 2024), NeedleBench (Li et al., 2024), LongBench (Bai et al., 2023), and Longcite.
5.2 Main Results
In Table 5, we compare our method HBP with Long-Align (Bai et al., 2024a) and packing method ISF (Yao et al., 2024). Our method achieves significant improvements in speed, outperforming Long-Align-packing, Long-Align-sorted, ISF by 2.0, 9.0, and 1.4. We also notice that LongAlign improves the average general tasks to some extent; it sacrifices performance in long tasks e.g., Ruler-128K. In contrast, our method maintains strong performance both in general short tasks and long tasks. The improvements are consistent across a wide range of model sizes, from 8B to 236B parameters. Notably, for the largest MOE model, DeepSeek-V2 (236B), our method achieves an impressive 2.4 training speed-up, reducing training time from 57.1 to 23.8 GPU days. Full results are shown in Appendix A.
5.3 Packing Strategy Results
Table 8 and Table 9 illustrate the complexity of different packing strategies and their corresponding final training times. It also shows that the training indicators Data Balance Ratio (DBR), Padding Ratio (PR), and Attention Balance Ratio (ABR) are strongly correlated with the training time, emphasizing the effectiveness of these indicators. Based on the computational complexity of packing strategies and their training time and accuracy performance, we ultimately selected ISF as our naive packing baseline.
| No | - | 0.64 | 0.744 | 0 | 57.6 | 42. | 16.7 (1.00) |
| Random | O(N) | 0 | 0.620 | 0.05 | 57.6 | 42.7 | 4.44 (3.76) |
| ISF | C*O(N+M) | 0 | 0.616 | 0.002 | 58.1 | 43.8 | 4.00 (4.17) |
| FFS | O(NM) | 0 | 0.618 | 0.001 | 8.3 | 42.1 | 4.02 (4.15) |
| FFD | O(NM) | 0 | 0.693 | 0.001 | 58.0 | 43.1 | 4.51 (3.70) |
| BFS | O(NM) | 0 | 0.618 | 0.001 | 57.5 | 42.1 | 4.11 (4.06) |
| SPFHP | O(N+S2) | 0 | 0.695 | 0.001 | 58.6 | 40.6 | 4.24 (3.94) |
| - | - | 0 | 0.878 | 0.676 | 48.2 | 46.4 | 11.4 (1.0) |
| random | O(N) | 0 | 0.648 | 0.089 | 51.44 | 47.8 | 3.64 (3.1) |
| ISF | C*O(N+M) | 0 | 0.64 | 0.022 | 50.9 | 47.8 | 3.28 (3.5) |
| FFS | O(NM) | 0 | 0.638 | 0.022 | 52.4 | 48.4 | 3.29 (3.5) |
| FFD | O(NM) | 0 | 0.71 | 0.022 | 50.6 | 48.4 | 3.45 (3.3) |
| BFS | O(NM) | 0 | 0.64 | 0.022 | 52.3 | 47.7 | 3.33 (3.4) |
| SPFHP | O(N+S2) | 0 | 0.71 | 0.029 | 52.0 | 47.8 | 3.47 (3.3) |
5.4 Ablation Results
Importance of Hybrid Training. Table 6 shows that both short-context and long-context data are important to maintain models’ capability in general tasks and long tasks. Lacking of long-context data (row 1) significantly impacts long-text capabilities. Lacking short-context data (row 2) sacrifices general abilities. These claims are further validated by only including partial short data in Longcite (row 3).
Components of HBP. In Table 7, we show that the Attention Balance Ratio (ABR) and Communication Ratio (CR) can be reduced significantly with hierarchical packing. In particular, ABR drops from 0.506 to 0.288, and the CR drops from 1.0 to 0.173. By batching data with a similar complexity of attention computation, we have much more balanced mini-batches with low ABR, from 0.288 to 0.002. Overall, we achieve a 1.4 speedup. Similar results can be observed in the larger LLaMA-3.1-70B model.
Curriculum Learning. In Table 10, we present the impact of curriculum learning on HBP. It is evident that starting with short general tasks and then transitioning to a mix of short- and long-context tasks is more beneficial for model training and convergence. We also designed a similar curriculum learning mechanism for the naive ISF baseline using a sophisticated sampling strategy, which shows improvements. This demonstrates that curriculum learning is a generalizable strategy for long-context SFT. Notably, since HBP naturally separates short and long contexts into different groups, employing curriculum learning becomes more straightforward and convenient.
| Model | CL | AVE | LongBench |
| LLama3.1-8B-HBP | 56.6 | 41.6 | |
| LLama3.1-8B-HBP | ✓ | 58.2 | 43.2 |
| LLama3.1-70B-HBP | 72.2 | 51.5 | |
| LLama3.1-70B-HBP | ✓ | 74.1 | 52.2 |
| LLama3.1-8B-ISF | 56.0 | 44.0 | |
| LLama3.1-8B-ISF | ✓ | 57.4 | 43.4 |
Stable Loss Normalizer. We compared several loss normalization methods by training models using different normalizers while keeping all other training configurations consistent. As shown in Table 11, the Ave-Token loss normalizer achieved the highest performance in both general tasks Average (AVE) 58.2 and long context tasks LongBench 43.1, Ruler-128K 70.8, and LongSite 71.5.
@@ 5.4 Ablation Results
| Ave-Token | 58.2 | 43.1 | 70.8 | 71.5 |
5.5 Importance of Hierarchical Groups Auto-Selection
Table 12 presents an example of 32K token-length training, showcasing various SP degrees and GC configurations. The second row highlights the minimal configuration that satisfies memory constraints. While this configuration adheres to memory requirements, it fails to achieve optimal performance. In contrast, the fourth row illustrates a more balanced and effective configuration with , achieving the fastest speed. These experiments demonstrate that, given a specific packing length, there are significant performance differences among various SP degrees and GC configurations. Table 13 presents the training speed of the optimal SP degrees and GC configurations for various sequence lengths . This indicates that the optimal training strategy is distinct for different packing groups. The evidence above emphasizes the importance of searching for the best groups and their corresponding training strategies, i.e., Auto-Group Selection.
| SP | GC Layer | AVE | Memory | Iter Time | GPU Days |
| 1 | 32 | - | OOM | - | - |
| 2 | 28 | 57.5 | 78G | 3.01 s | 3.73 |
| 4 | 23 | 57.8 | 78G | 2.95 s | 3.37 |
| 8 | 8 | 57.6 | 77G | 2.82 s | 3.04 |
| 16 | 0 | 56.6 | 76G | 3.60 s | 3.93 |
| Packing Groups | SP | GC Layer | Memory | Iter Time |
| 8k | 2 | 8 | 77 G | 2.69 s |
| 16k | 1 | 28 | 76 G | 2.65 s |
| 32k | 8 | 8 | 76 G | 2.83 s |
| 64k | 4 | 28 | 78 G | 3.01 s |
| 128k | 8 | 28 | 79 G | 3.05 s |
| 0.513 | 52.1 | 47.2 | 87.7 | 72.1 | 2.95 (1.0) | ||
| ✓ | 0.269 | 53 | 47.1 | 89.1 | 72.0 | 2.33 (1.27) | |
| ✓ | ✓ | 0.004 | 52.4 | 47.1 | 88.3 | 72.3 | 2.04 (1.44) |
| Model | Dataset | AVe | LongBench | LongWriter | GPU Days (speed up) |
| Llama3.1-8B-ISF | Tulu3 + LongWriter | 57.3 | 40.9 | 67.0 | 4.4 (1.0) |
| Llama3.1-8B-HBP | Tulu3 + LongWriter | 57.8 | 42.7 | 66.4 | 3.5 (1.26) |
5.6 Dataset Generalization
5.6.1 OpenHermes
5.6.2 LongWriter
@@ 6 Conclusion and Limitations
In this paper, we proposed Hierarchical Balance Packing (HBP), a novel strategy to address workload imbalances in long-context LLM training through multi-level data packing and a dynamic training pipeline. Due to constraints in computational resources and open-source training datasets, we have not tested on longer contexts, such as 256K or 512K. Additionally, we have not validated HBP on other post-training tasks, such as RLHF or DPO. We leave these explorations to future work.
Appendix A Full Results
Tables 16 and 17 present our complete results for general tasks and long-context tasks, respectively. These results collectively validate the effectiveness of our HBP method.

### v3

@@ 5.2 Main Results
[…]
In Table 5, we compare our method HBP with LongAlign [6] and packing method ISF [8]. We also notice that LongAlign improves the average general tasks to some extent; it sacrifices performance in long tasks e.g., Ruler-128K. In contrast, our method maintains strong performance both in general short tasks and long tasks. The improvements are consistent across a wide range of model sizes, from 8B to 236B parameters. Notably, for the largest MoE model, DeepSeek-V2 (236B), our method achieves an impressive 2.4 training speed-up, reducing training time from 57.1 to 23.8 GPU days. Full results are shown in Appendix C.
| Model | Hierarchical | Balance | ABR | CR | AVE | GPU Days (speed up) |
| LLaMA3.1-8B | 0.506 | 1.0 | 56.0 | 5.22 (1.0) | ||
| LLaMA3.1-8B | ✓ | 0.288 | 0.173 | 56.4 | 4.51 (1.2) | |
[…]
| LLaMA3.1-70B | 0.506 | 1.0 | 72.2 | 44.4 (1.0) | ||
| LLaMA3.1-70B | ✓ | 0.288 | 0.173 | 72.8 | 33.3 (1.25) | |
[…]
5.3 Ablation Results
Importance of Hybrid Training. Table 10 shows that both short- and long-context data are essential for maintaining general and long-text capabilities. "Longcite (8K)" includes only sequences up to 8K. All settings use naive packing. Removing long-context data (row 1) harms long-text performance, while removing short-context data (row 2) weakens the general ability. Row 3, with partial short data, confirms these effects.
| Model | CL | AVE | LongBench |
| LLama3.1-8B-HBP | 56.6 | 41.6 | |
[…]
| LLama3.1-70B-HBP | 72.2 | 51.5 | |
[…]
| LLama3.1-8B-ISF | 56.0 | 44.0 | |
[…]
| Model | CL-Iterations | AVE | LongBench |
| LLama3.1-8B-HBP | 0 | 56.6 | 41.6 |
| LLama3.1-8B-HBP | 100 | 58.0 | 43.1 |
[…]
| LLama3.1-70B-HBP | 0 | 72.2 | 51.5 |
| LLama3.1-70B-HBP | 100 | 73.8 | 52.4 |
[…]
Components of HBP. In Table 6, we show that the Attention Balance Ratio (ABR) and Communication Ratio (CR) can be reduced significantly with hierarchical packing. In particular, ABR drops from 0.506 to 0.288, and the CR drops from 1.0 to 0.173. By balancing the batching data with a similar complexity of attention computation, we have much more balanced mini-batches with low ABR, from 0.288 to 0.002. Overall, we achieve a 1.4 speedup. Similar results can be observed in the larger LLaMA-3.1-70B model.
Curriculum Learning. Table 8 presents the impact of curriculum learning on HBP. Starting with short tasks and gradually mixing in long-context tasks benefits training and convergence. We applied a similar curriculum strategy to the naive ISF baseline using advanced sampling, which also showed improvements. This confirms the general effectiveness of curriculum learning for long-context SFT. Notably, HBP naturally separates short and long contexts, making curriculum learning easier to apply.
Iterations for Curriculum Learning training. Table 8 presents detailed ablation studies across different models. As shown, even a simple curriculum learning setup, adding 100 steps of early-stage training with short samples, already yields noticeable improvements. To ensure training stability, we ultimately adopt a setting with 500 steps of short-sample-only training in our final experiments.
@@ 5.3 Ablation Results
[…]
@@
[…]
Stable Loss Normalizer. We compared several loss normalization methods by training models using different normalizers while keeping all other training configurations consistent. As shown in Table 10, the Ave-Token loss normalizer achieved the highest performance in both general tasks, Average (AVE) 57.6, and long context tasks LongBench 43.1, Ruler-128K 70.8, and LongSite 71.2.
Importance of Hierarchical Groups Auto-Selection. Table 11 compares manual (row 2) and automatic (row 3) packing groups. Automatic selection of groups (based on Appendix F) achieves greater speedup, reducing training time from 4.25 to 3.73 GPU days.
| Model | Groups | Selection | SP | ABR | CR | AVE | LongBench | GPU Days (speed up) |
| LLama3.1-8B-ISF | [128K] | manual | [8] | 0.506 | 1.0 | 56.0 | 44.0 | 5.22 (1.0) |
| LLama3.1-8B-HBP | [32K, 128K] | manual | [2, 8] | 0.002 | 1.0 | 57.9 | 43.2 | 4.25 (1.22) |
[…]
Compared with FlexSP. As shown in Table 12, HBP achieves lower ABR, DBR, and PR than FlexSP, while maintaining a similar CR, contributing to faster training (4.35 to 3.73 GPU days), even considering only the model computation time. Furthermore, FlexSP introduces a per-batch grouping overhead (5-15 seconds), which further slows down training (see Appendix B). HBP performs global data-level operations in advance with negligible data overhead.
Negligible Overhead. The overhead of HBP is negligible: Auto-Selection takes minutes and Packing seconds, together contributing less than of training time (see Appendix G).
| Model | SP | CL | PR | DBR | CR | ABR | AVE | LongBench | GPU Days (speed up) |
| LLama3.1-8B-ISF | [8] | 0 | 0.001 | 1.0 | 0.506 | 56.0 | 44.0 | 5.22 (1.0) | |
| LLama3.1-8B-FlexSP [24] | [1,2,4,8] | 0.04 | 0.064 | 0.171 | 0.36 | 56.1 | 41.8 | 4.35 (1.2) | |
| LLama3.1-8B-HBP | [1,8] | 0 | 0.001 | 0.173 | 0.002 | 56.6 | 41.6 | 3.73 (1.4) | |
[…]
@@ 6 Conclusion and Limitations
In this paper, we proposed Hierarchical Balance Packing (HBP), a novel strategy to address workload imbalances in long-context LLM training through multi-level data packing and a dynamic training pipeline. Due to limitations in computational resources, we have not conducted experiments on longer contexts, such as 256K or 512K. Additionally, we have not validated HBP on other post-training tasks, such as RLHF or DPO. We leave these explorations to future work.
Acknowledgements
This work was supported by the National Key Research and Development Program of China (Grant No. 2023YFB4405102), the Postdoctoral Fellowship Program and the China Postdoctoral Science Foundation (Grant No. BX20250487), and in part by the Natural Science Foundation of Hunan Province (Grant No. 2024JJ6525).
NeurIPS Paper Checklist
1.
Claims
Question: Do the main claims made in the abstract and introduction accurately reflect the contributions and scope of the paper?
Answer: [Yes]
Justification: The main claims in the abstract and introduction accurately reflect the paper’s contributions and scope.
Guidelines:
The answer NA means that the abstract and introduction do not include the claims made in the paper.
The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
2.
Limitations
Question: Does the paper discuss the limitations of the work performed by the authors?
Answer: [Yes]
Justification: The paper provides a brief discussion of the limitations in the Conclusion section.
Guidelines:
The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
The authors are encouraged to create a separate "Limitations" section in their paper.
The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
3.
Theory assumptions and proofs
Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
Answer: [N/A]
Justification: The paper is primarily based on empirical observations and proposes heuristic approaches, which are not fully supported by formal mathematical proofs or a complete set of theoretical assumptions.
Guidelines:
The answer NA means that the paper does not include theoretical results.
All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
All assumptions should be clearly stated or referenced in the statement of any theorems.
The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
Theorems and Lemmas that the proof relies upon should be properly referenced.
4.
Experimental result reproducibility
Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
Answer: [Yes]
Justification: The experimental results can be fully reproduced based on the information provided in the paper, and the authors have indicated that the code will be released in the future.
Guidelines:
The answer NA means that the paper does not include experiments.
If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
5.
Open access to data and code
Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
Answer: [N/A]
Justification: The code and data are not yet publicly available, and although the authors plan to release them later, they are currently not provided with sufficient instructions in the supplemental material to ensure faithful reproduction of the results.
Guidelines:
The answer NA means that paper does not include experiments requiring code.
Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
6.
Experimental setting/details
Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
Answer: [Yes]
Justification: The full details are provided in the main text and supplemental material.
Guidelines:
The answer NA means that the paper does not include experiments.
The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
The full details can be provided either with the code, in appendix, or as supplemental material.
7.
Experiment statistical significance
Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
Answer: [Yes]
Justification: The results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
Guidelines:
The answer NA means that the paper does not include experiments.
The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
The assumptions made should be given (e.g., Normally distributed errors).
It should be clear whether the error bar is the standard deviation or the standard error of the mean.
It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
8.
Experiments compute resources
Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
Answer: [Yes]
Justification: The paper provides sufficient information about the computational resources used, including details in both the main text and the Appendix.
Guidelines:
The answer NA means that the paper does not include experiments.
The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
9.
Code of ethics
Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
Answer: [Yes]
Justification: The research presented in the paper fully conforms to the NeurIPS Code of Ethics in all respects.
Guidelines:
The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
10.
Broader impacts
Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
Answer: [N/A]
Justification: There is no societal impact of the work performed.
Guidelines:
The answer NA means that there is no societal impact of the work performed.
If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
11.
Safeguards
Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
Answer: [N/A]
Justification: The paper poses no such risks.
Guidelines:
The answer NA means that the paper poses no such risks.
Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
12.
Licenses for existing assets
Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
Answer: [Yes]
Justification: All resources used in the paper comply with their respective licenses and terms of use.
Guidelines:
The answer NA means that the paper does not use existing assets.
The authors should cite the original paper that produced the code package or dataset.
The authors should state which version of the asset is used and, if possible, include a URL.
The name of the license (e.g., CC-BY 4.0) should be included for each asset.
For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
13.
New assets
Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
Answer: [N/A]
Justification: The paper does not release new assets.
Guidelines:
The answer NA means that the paper does not release new assets.
Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
The paper should discuss whether and how consent was obtained from people whose asset is used.
At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
14.
Crowdsourcing and research with human subjects
Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
Answer: [N/A]
Justification: The paper does not involve crowdsourcing nor research with human subjects.
Guidelines:
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
15.
Institutional review board (IRB) approvals or equivalent for research with human subjects
Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
Answer: [N/A]
Justification: The paper does not involve crowdsourcing nor research with human subjects.
Guidelines:
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
16.
Declaration of LLM usage
Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
Answer: [N/A]
Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
Guidelines:
The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Packing Strategy Results
Table 13 illustrates the complexity of different packing strategies and their corresponding final training times. It also shows that the training indicators Data Balance Ratio (DBR), Padding Ratio (PR), and Attention Balance Ratio (ABR) are strongly correlated with the training time, emphasizing the effectiveness of these indicators. Based on the computational complexity of packing strategies and their training time, and accuracy performance, we ultimately selected ISF as our naive packing baseline.
| No | - | 0.63 | 0.72 | 0 | 56.5 | 41.8 | 38.3 (1.00) |
| Random | O(N) | 0 | 0.53 | 0.03 | 56.2 | 42.0 | 5.57 (6.87) |
[…]
| FFS | O(NM) | 0 | 0.508 | 0.01 | 56.3 | 42.0 | 5.24 (7.30) |
| FFD | O(NM) | 0 | 0.515 | 0.01 | 56.4 | 42.4 | 5.48 (6.98) |
| BFS | O(NM) | 0 | 0.508 | 0.01 | 56.8 | 41.7 | 5.24 (7.30) |
| SPFHP | O(N+S2) | 0 | 0.512 | 0.01 | 56.2 | 42.3 | 5.40 (7.10) |
Appendix B Results of FlexSP data grouping
The experiments were conducted with a sequence length of 128K using 32 GPUs. The results for FlexSP were obtained by directly testing with the officially released code. The FlexSP data group method incurs a non-negligible overhead of 5–15 seconds per batch, which cannot always be hidden by the training time shown in Table 14 Rows 3 and 4.
| Model | Dataset | Group method | Data Time | Model Time | Total Time |
| GPT-7b | github | FlexSP | 0.25s | 15.4s | 15.6s |
| GPT-7b | CommonCrawl | FlexSP | 0.3s | 13.0s | 13.3s |
| GPT-7b | Wikipedia | FlexSP | 4.41s | 2.4s | 6.81s |
| GPT-7b | Tulu3 + Longcite | FlexSP | 3.43s | 3.02s | 6.45s |
[…]
Appendix C Full Results
Tables 15 and 16 present our complete results for general tasks and long-context tasks, respectively. These results collectively validate the effectiveness of our HBP method. The other dense models all use the GQA architecture, while Deepseek-V2 adopts the MLA structure, which involves more communication due to a larger number of K and V heads, similar to MHA. HBP effectively reduces the communication overhead, making the speedup more significant.